# Evaluating LLMs’ Mathematical Reasoning in Financial Document Question Answering

### v2

@@ Abstract
Large Language Models (LLMs), excel in natural language understanding, but their capability for complex mathematical reasoning with an amalgamation of structured tables and unstructured text is uncertain. This study explores LLMs’ mathematical reasoning on four financial tabular question-answering datasets: TATQA, FinQA, ConvFinQA, and Multihiertt. Through extensive experiments with various models and prompting techniques, we assess how LLMs adapt to complex tables and mathematical tasks. We focus on sensitivity to table complexity and performance variations with an increasing number of arithmetic reasoning steps. The results provide insights into LLMs’ capabilities and limitations in handling complex mathematical scenarios for semi-structured tables. Ultimately, we introduce a novel prompting technique tailored to semi-structured documents, matching or outperforming other baselines in performance while providing a nuanced understanding of LLMs abilities for such a task.
@@ 1 Introduction
In the constantly evolving realm of artificial intelligence, Large Language Models (LLMs) have risen as cutting-edge tools for natural language understanding. They excel in a wide array of NLP tasks, including machine translation (MT), text summarization, question answering, and code generation. One specific area where LLMs’ mathematical reasoning abilities come under scrutiny is the domain of numerical reasoning tasks. Past research has delved into the potential of language models for mathematical reasoning tasks, as seen in studies such as Amini et al. (2019); Upadhyay and Chang (2017); Patel et al. (2021); Cobbe et al. (2021). These investigations provide a means to evaluate the performance of language models when it comes to solving mathematical problems, ranging from straightforward math word problems to more complex ones.
Figure 1: An example of a QA pair on a financial table.
@@
However, the problem becomes significantly more challenging when we encounter a hybrid of structured and unstructured data, as exemplified by semi-structured tables. Such tables are common in documents like invoices, health records, and financial reports in corporate settings. In the financial domain, these tables present numerical data in a structured format, including income statements, balance sheets, cash flow statements, shareholder equity data, and annual reports. The majority of NLP models are primarily trained to handle unstructured textual data, which limits their ability to reason over semi-structured data, such as tables, or more intricate hybrids of tables and text, as seen in Chen et al. (2020b); Aly et al. (2021); Chen et al. (2020a, 2021a). Tables, especially those with intricate relationships and dependencies, often necessitate multi-hop reasoning, connecting information across multiple steps.
@@
NLP models may encounter difficulties in performing such multi-step reasoning, particularly when dealing with complex mathematical operations involving tables, as highlighted in Li et al. (2022). Previous research Chen (2023), exemplified these issues and demonstrated LLM capacity to process and reason with semi-structured tables. However, these studies are somewhat constrained and don’t explicitly explore the models’ mathematical reasoning abilities. This is particularly evident in data/tasks that involve a substantial number of arithmetic reasoning steps, operate on extreme orders of magnitude, or deal with intricate tables where extracting the relevant information for a query becomes challenging.
@@
Moreover, when handling domain-specific documents, such as those in finance, a language model must not only have the necessary domain knowledge to craft the right approach for task-solving but also the capability to manipulate structured data, such as tables. Therefore, in this study, we aim to fill this gap by providing both qualitative and quantitative analyses of LLM’s ability to reason on mathematical content on four finance datasets. Our goal is to illustrate how model performance varies as the numerical complexity of the underlying data and the intricacy of mathematical reasoning steps required to solve a query increase. Our contributions are as follows:
@@
1. We conduct a comprehensive evaluation of state-of-the-art Large Language Models (LLMs) in tabular question answering, with a specific focus on mathematical reasoning tasks, using public financial tabular datasets to establish a robust performance benchmark.
@@
2. Our analysis is thorough and multifaceted, encompassing both qualitative and quantitative aspects. We aim to provide nuanced insights into the strengths and limitations of LLMs in tabular question answering, especially in scenarios involving mathematical reasoning.
3. Building upon qualitative analysis, we introduce a novel and improved prompting method called EEDP (Elicit $\longrightarrow$ Extract $\longrightarrow$ Decompose $\longrightarrow$ Predict). This innovative approach yields comparable or enhanced performance when compared to existing prompt engineering methods. This method of explicitly prompting the model to break it’s response into sections provides a more explainable and nuanced understanding of their weaknesses in performing such tasks.
2 Related Works
2.1 LLMs on Mathematical Reasoning
2.2 Tabular Question Answering
Handling diverse input formats in question answering, including structured tables and visual data, poses challenges for language models. HybridQA Chen et al. (2020b) introduces questions requiring reasoning over tables and text. MultimodalQA Talmor et al. (2021) adds visual inference. Our focus is on multi-hop question answering over tables and text. TAPAS Herzig et al. (2020) tackles table-based questions without logical forms, while Tapex Liu et al. (2022) empowers generative models with table reasoning. Models like FinQANet, TagOP, and MT2Net Chen et al. (2021b); Zhu et al. (2021); Zhao et al. (2022) consist of a retriever module to first retrieve the supporting facts from the input financial report followed by a reasoning module to perform reasoning over the retrieved facts and get the final answer. Both Retrieving the right evidences from financial reports and reasoning over the retrieved facts require in-domain knowledge. For instance, questions may require understanding the financial implications of transactions, identifying trends in revenue or interpreting complex statements regarding investments. This necessitates a reliance on an external knowledge base to overcome the challenges faced while dealing with domain-specific multi-hop question answering.
2.3 Prompt Engineering
In-context Learning (ICL) equips Language Models (LLMs) with task examples and queries, enabling them to perform target tasks without updating model parameters (Brown et al., 2020; OpenAI, 2023). They excel in mathematical reasoning with few-shot prompts but struggle with more complex tasks. Methods like chain-of-thoughts (CoT) (Wei et al., 2022) have improved LLM performance by guiding them through intermediate reasoning steps. Enhancing multi-step reasoning involves two main approaches: improving in-context examples and obtaining better reasoning steps. Some focus on more stable in-context example selection (Fu et al., 2023; Rubin et al., 2022; Lu et al., 2023). Others adopt a modular approach, using off-the-shelf tools (Schick et al., 2023), program of thoughts (PoT) (Chen et al., 2022a), or decomposition strategies (Zhou et al., 2023; Dua et al., 2022; Khot et al., 2023). Our task requires complex multi-step reasoning across diverse information sources. LLMs, as demonstrated by (Chen, 2023), can reason over structured tables without explicit encoding. They also serve as versatile decomposers, breaking down extensive evidence and complex questions (Ye et al., 2023b).
3 Dataset Annotations
In our evaluations, we considered four tabular datasets: FinQA Chen et al. (2021b), TATQA Zhu et al. (2021), ConvFinQA Chen et al. (2022b), and Multihiertt Zhao et al. (2022). These datasets involve questions requiring numerical reasoning over semi-structured data. We excluded simple span selection questions from TATQA. Each dataset provides pre-annotated executable programs for deriving ground-truth answers. We meticulously annotate the datasets considered for our study with the following metadata:
@@ 1. Number of Reasoning Steps:
Including the count of arithmetic operations in questions is crucial. More operations reflect increased complexity in reasoning, and their interdependence offers insights into the models’ proficiency. This annotation, applied across all four datasets, reveals their ability in handling intricate arithmetic tasks.
Refer to Figure 5 in Appendix (A.2), answering the question require multiple reasoning steps.
@@ 2. Question Categorization:
In numerical reasoning, grasping the evolution from fundamental arithmetic to advanced operations is crucial, marking a shift in cognitive complexity. As questions advance, they typically involve more intricate combinations of operations and linguistic nuances. Our research seeks to identify both the capabilities and limitations of LLMs in understanding these concepts.
We establish 12 mathematical concept categories (Table 1) with corresponding definitions, annotating each question. The dataset coverage across these categories is shown in Figure 6 in Appendix (A.2). Notably, categories like Division and Ratio share similarities but differ in focus: Division involves the division operator, while Ratio encompasses ratios, fractions, and inverse problems. Change in Ratio questions add complexity with quantity changes requiring subtraction. Additionally, we omit NEED-IN-DOMAIN-INFO due to domain-specific knowledge focus and TIME questions due to limited sample size.
@@
Table 1: Mathematical Concept Categories and Definitions for Studying Concept Comprehension Abilities of LLMs
3. Number of Rows:
Evaluating performance with larger supporting tables is crucial. Larger tables complicate multi-hop reasoning tasks by increasing the amount of information, making it harder to identify relevant evidence. We prioritize these annotations for datasets like FinQA and Multihiertt, where questions mainly use tables as supporting evidence. Therefore, these annotations are confined to these datasets. In Multihiertt, when multiple tables support evidence, we consider the one with the highest row count. The dataset coverage across the number of rows for Multihiertt and FinQA is shown in Figure 7 in Appendix (A.2).
4. Depth of Hierarchy:
In hierarchical tables, such as those in Multihiertt, evaluating model performance concerning the growing depth of hierarchy in cells with critical information becomes paramount. To tackle this, we annotate each example in Multihiertt with the hierarchy depth of cells containing crucial information, considering the maximum hierarchy value for multiple value extraction. Our approach to estimating hierarchy depth is illustrated in Figure 16, and Figure 8 in Appendix (A.2) displays the sample distribution that require the extraction of gold evidence from various hierarchy depths.
5. Proportion of Empty Cells:
@@
Assessing empty cell proportions is crucial to quantify data ambiguity. More empty cells suggest higher uncertainty, which can hinder models’ ability to derive meaningful insights and impact reasoning accuracy. In Multihiertt, where tables are hierarchical in nature and empty cells occur quite frequentlt, we annotate examples with the percentage of empty cells, contributing to our understanding of data ambiguity. For the distribution of empty cells, refer to Figure 9 in Appendix (A.2).
4 Model Performance
Models:
@@ 4 Model Performance
In this study, we choose to experiment with LLMs like GPT-3.5-Turbo, GPT-4, PaLM-540B, Mistral-7B-Instruct [^1] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, Llama-2-13B [^2] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf and MAmmoTH-13B [^3] https://huggingface.co/TIGER-Lab/MAmmoTH-13B.
Baseline Prompting Strategies:
@@
Prompt Engineering involves designing inputs for LLMs to produce optimal results. Typically, for an Instruction-tuned LLM, it’s assumed that we give the model a task-specific instruction $\mathcal{I}$ accompanied with a few (usually $k\in\{2,4\}$) demonstrations $\mathcal{D_{T}}$ of a task T. We experiment with the following prompting techniques as baselines.
@@
  • Direct: In this setup, we explicitly instruct the models to abstain from providing explanations and just return the final answer to the posed question. For this scenario, $\mathcal{D_{T}}$ contains $\{(p_{i},q_{i},a_{i})\}_{i=1}^{k}$ where p is the premise which includes the table and the associated text, q is the question and a is the ground-truth answer.
@@
  • CoT: Moving forward, we experiment with the chain-of-thoughts prompting strategy where we instruct the model to output the explanation to the answer derived by it. Here, our $\mathcal{D_{T}}$ contains $\{(p_{i},q_{i},e_{i})\}_{i=1}^{k}$ where p is the premise which includes the table and the associated text, q is the question and e is the explanation of the answer.
@@
  • PoT: In this case, the expected response is a code derivation of the final answer. Here, our $\mathcal{D_{T}}$ contains $\{(p_{i},q_{i},c_{i})\}_{i=1}^{k}$ where p is the premise which includes the table and the associated text, q is the question and c is the code-derivation of the answer.
@@
  • Decomposers: Ye et al. (2023a) proposed to address the challenge of handling large tables by decomposing them into more manageable subtables. Similarly, we handle complex questions by breaking them down into simpler subquestions. In their work, they show that their method works well on SQL tables where the table can easily be shortened to remove distractive information without the loss of any supporting evidence. Next, they parse the question to decompose it into simpler, more manageable subquestions. Finally, the model is expected to answer each of the subquestions and compose those to arrive at the final answer. In this case, our demonstration set $\mathcal{D_{T}}$ contains $\{(p_{i}^{{}^{\prime}},\langle q_{1},q_{2},...,q_{n}\rangle,a_{i})\}_{i=1}^{k}$ where $p^{{}^{\prime}}$ is the premise obtained by the removal of information irrelevant to the question from the original premise p and $\langle q_{1},q_{2},...,q_{n}\rangle$ are the subquestions whose answer lead to the final answer.
EEDP (Ours):
@@
We propose a novel prompting strategy: Elicit $\longrightarrow$ Extract $\longrightarrow$ Decompose $\longrightarrow$ Predict.
@@
    Predict: The model finally returns the derived answer in the above steps,
Comparison with other state-of-the-art methods
  1. 1.
    PoT vs EEDP: PoT enhances LLM inference with the use of variable names for the supporting values extracted from the premise and prompts the LLM to express their thought process in the form of programs. The model output is a program which is executed externally to derive the final answer. EEDP proposes to decompose a complex reasoning task into simple atomic steps whose solutions can be composed to give the final answer. In PoT, they don’t make the language model do the computation while in our case the language model not only outputs the reason but also computes the final answer. This distinction implies that PoT may have an inherent advantage over EEDP.
  2. 2.
    Decomposers vs EEDP: The prompting strategy proposed by Ye et al. (2023b) was originally designed for quering SQL tables, they use LLMs to break down evidence and questions for SQL interpreters. In contrast, our approach addresses more complex scenarios involving both tables and text, requiring advanced reasoning skills beyond the capability of standard SQL interpreters. Pruning a non-SQL table using this method can lead to significant information loss from the premise which can be a potential ingredient required to derive the final answer. Additionally, this is an expensive method as it requires 3X API calls as opposed to other prompting methods. Moreover, EEDP is a unified prompting strategy which integrates multiple solver elements into a single unified prompt for elicitation, extraction, decomposition and prediction.
Table 2. The reason behind this was statistically found to be inaccurate formation of subtables from the main table leading to information loss as described in the previous paragraph. The performance of EEDP either surpasses or matches very closely with that of PoT. The number of shots were adjusted depending upon the context length of the model. We can see that MammoTH-13B which is finetuned on the MathInstruct dataset Yue et al. (2023) containing Instruction-Response pairs where the responses are a hybrid of CoT and PoT rationales fails to do well on this task with EEDP methodology. We argue that this is due to two potential reasons:
  1. 1.
    Reduction of the number of shots to adjust the context length as the EEDP response is longer than that of the other methods.
  2. 2.
    Finetuning may contribute to suboptimal performance due to its alignment with a particular style and format of responses, potentially limiting the model’s adaptability and generalization to diverse contexts.
Fig. 17 shows a template for EEDP strategy. Table 2 shows a comparison in performance between different prompting strategies.
Model Dataset Direct CoT PoT EEDP Decomposers
TATQA 55.81 86.91 89.99 88.67 47.46
GPT-4 FinQA 65.12 72.38 75.26 76.05 44.93
ConvFinQA 63.10 71.19 78.81 77.91 18.76
Multihiertt 41.35 55.13 67.23 70.32 36.86
TATQA 31.38 77.57 82.11 79.73 28.53
GPT-3.5-Turbo FinQA 40.47 59.18 68.97 61.88 32.33
ConvFinQA 37.62 48.33 61.19 61.75 10.50
Multihiertt 25.88 42.33 52.18 49.65 20.61
TATQA 44.66 62.93 61.60 81.51 57.94
PaLM 2-540B FinQA 30.33 34.79 30.41 61.95 46.38
ConvFinQA 20.19 38.00 40.14 63.42 22.32
Multihiertt 14.20 20.67 36.52 37.97 20.19
TATQA 3.36 35.95 34.16 40.95 25.93
Llama 2-13B FinQA 1.80 25.34 12.97 30.47 11.91
ConvFinQA 3.80 29.45 29.92 39.42 10.35
Multihiertt 1.54 30.66 18.12 24.15 16.86
TATQA 19.11 56.25 10.02 4.37 22.89
MammoTH-13B FinQA 22.83 35.32 15.86 35.05 17.65
ConvFinQA 21.61 46.08 8.78 32.77 7.83
Multihiertt 10.12 18.56 6.57 18.36 11.87
TATQA 10.92 59.14 16.53 56.06 7.24
Mistral-7B FinQA 26.11 34.23 10.56 34.86 12.34
ConvFinQA 12.35 48.45 14.48 36.57 11.16
Multihiertt 14.909 22.92 14.94 10.97 11.63
@@
RQ: Where do LLMs falter while performing hybrid question answering involving arithmetic operations?
We categorize errors and identify their origins through a detailed analysis of model responses.

### v3

@@ Abstract
Large Language Models (LLMs), excel in natural language understanding, but their capability for complex mathematical reasoning with a hybrid of structured tables and unstructured text remain uncertain. This study explores LLMs’ mathematical reasoning on four financial tabular question-answering datasets: TATQA, FinQA, ConvFinQA, and Multihiertt. Through extensive experiments with various models and prompting techniques, we assess how LLMs adapt to complex tables and mathematical tasks. We focus on sensitivity to table complexity and performance variations with an increasing number of arithmetic reasoning steps. The results provide insights into LLMs’ capabilities and limitations in handling complex mathematical scenarios for semi-structured tables. Ultimately, we introduce a novel prompting technique EEDP tailored to semi-structured documents, matching or outperforming baselines performance while providing a nuanced understanding of LLMs abilities.
@@ 1 Introduction
In the constantly evolving realm of artificial intelligence, Large Language Models (LLMs) have risen as cutting-edge tools for natural language understanding. They excel in a wide array of NLP tasks, including machine translation (MT), text summarization, question answering, and code generation. One specific area where LLMs’ mathematical reasoning abilities come under scrutiny is the domain of numerical reasoning tasks. Past research has delved into the potential of language models for mathematical reasoning tasks, as seen in studies such as in Amini et al. (2019); Upadhyay and Chang (2017); Patel et al. (2021); Cobbe et al. (2021). These investigations provide a means to evaluate the performance of language models when it comes to solving mathematical problems, ranging from straightforward math word problems to more complex ones.
The Goldman Sachs Group Incorporation Notes to Consolidated Financial Statements
The table below presents a summary of Level 3 financial assets.
Financial Asset Dec. 2017 Cash Instruments \$15,395 Derivatives \$3,802 Other Financial Instruments \$4 Q: What was the total value of Level 3 financial assets for Goldman Sachs in December 2017? A: \$15,395 + \$3,802 + \$4 = \$19,201
Figure 1: An example of a semi-structured financial document question answering.
@@
However, the problem becomes significantly more challenging when we encounter a hybrid of structured such as semi-structured tables and unstructured text, as shown in example in figure 1. Such tables are common in documents such as invoices, health records, and financial reports in corporate settings. In the financial domain, these tables present numerical data in a structured format, including income statements, balance sheets, cash flow statements, shareholder equity data, and annual reports. The majority of NLP models are primarily trained to handle raw unstructured textual data, which limits their ability to reason over semi-structured data, such as tables, or more intricate hybrids of tables and text, as seen in Chen et al. (2020b); Aly et al. (2021); Chen et al. (2020a, 2021a). Tables, especially these with intricate relationships and dependencies, often necessitate multi-hop reasoning, connecting information across multiple steps, as shown in Gupta et al. (2020).
@@
NLP models may encounter difficulties in performing such multi-step reasoning, particularly when dealing with complex mathematical operations involving tables, as highlighted in Li et al. (2022). Previous research such as Chen (2023), exemplified these issues and demonstrated LLM capacity to process and reason with semi-structured tables. However, these studies are somewhat constrained and don’t explicitly explore the models’ mathematical reasoning abilities. This is particularly evident in data/tasks that involve a substantial number of arithmetic reasoning steps, operate on extreme orders of magnitude, or deal with intricate tables where extracting the relevant information for a query becomes challenging.
@@
Moreover, when handling domain-specific documents, such as those in finance, a language model must not only have the necessary domain knowledge to craft the right approach for task-solving but also the capability to manipulate structured data, such as tables. Therefore, in this study, we aim to fill this gap by providing both qualitative and quantitative analyses of LLM’s ability to reason on mathematical content on four finance datasets i.e. FinQA Chen et al. (2021b), TATQA Zhu et al. (2021), ConvFinQA Chen et al. (2022b), and Multihiertt Zhao et al. (2022). These datasets feature questions demanding intricate numerical reasoning, combining semi-structured tables and text. Each dataset provides pre-annotated executable programs for precise answer retrieval. Our goal is to illustrate how model performance varies as the numerical complexity of the underlying data and the intricacy of the mathematical reasoning steps required to solve a query increase. Building upon these observations, we propose a novel approach termed (Elicit $\longrightarrow$ Extract $\longrightarrow$ Decompose $\longrightarrow$ Predict) EEDP, designed to deconstruct model responses into discrete components. This innovative method offers a deeper, more transparent insight into the numerical limitations of the model when tackling these tasks. Our contributions are as follows:
@@
1. We conduct a comprehensive robust evaluation of state-of-the-art Large Language Models (LLMs) for tabular (hybrid) question answering, with a specific focus on mathematical reasoning tasks, using public financial tabular datasets to establish a thorough performance benchmark.
@@
2. Our analysis is thorough and multifaceted, encompassing both qualitative and quantitative aspects across several dimensions. We aim to provide nuanced insights into the strengths and limitations of LLMs in tabular (hybrid) question answering, especially in scenarios involving mathematical reasoning.
3. Building upon qualitative analysis, we introduce a novel and improved prompting method called EEDP. Our novel approach not only enhances our understanding of model weaknesses but also substantially enhances model performance compared to existing prompting methods across multiple models types.
Our metadata dataset and source code are available at https://vgupta123.github.io/eedp.
2 Metadata Annotations
We annotated four tabular datasets: FinQA, TATQA, ConvFinQA, and Multihiertt with meta information related to a.) reasoning steps, b.) question category, c.) table length, d.) hierarchical complexity e.) missing information.  [^1] One author annotated the data, and the other checked for accuracy; we took stringent measures to minimize errors. Below, we provide detailed information about these meta-data annotations:
@@ 1. Number of Reasoning Steps:
Including the count of arithmetic operations in questions is crucial. More operations reflect increased complexity in reasoning, and their interdependence offers insights into the models’ proficiency. This annotation, applied across all four datasets, reveals their ability in handling intricate arithmetic tasks. Refer to Figure 8 in Appendix A.4 for distribution of questions based on the number of reasoning steps involved.
@@ 2. Question Categorization:
In numerical reasoning, grasping the evolution from fundamental arithmetic to advanced operations is crucial, marking a shift in cognitive complexity. As questions advance, they typically involve more intricate combinations of operations and linguistic nuances. Our research identify both the capabilities and limitations of LLMs in understanding these concepts.
We establish 12 mathematical concept categories (Table 1) with corresponding definitions, annotating each question. The dataset coverage across these categories is shown in Figure 9 in Appendix A.4. Notably, categories like Division and Ratio share similarities but differ in focus: Division involves the division operator, while Ratio encompasses ratios, fractions, and inverse problems. Change in Ratio questions add complexity with quantity changes requiring subtraction. Additionally, we omit NEED-IN-DOMAIN-INFO due to domain-specific knowledge focus and TIME questions due to limited sample size.
@@
Table 1: Mathematical concept categories and definitions for studying LLM concept comprehension abilities.
3. Table Length:
Evaluating performance with larger supporting tables is crucial. Larger tables complicate multi-hop reasoning tasks by increasing the amount of information, making it harder to identify relevant evidence. We prioritize these annotations for datasets like FinQA and Multihiertt, where questions mainly use tables as supporting evidence. Therefore, these annotations are confined to these datasets. In Multihiertt, when multiple tables support evidence, we consider the one with the highest row count i.e. maximum table length. The dataset distribution for Multihiertt and FinQA w.r.t table length (number of rows) is shown in Figure 6 in Appendix A.4.
4. Hierarchical Complexity:
In hierarchical tables, such as those in Multihiertt, evaluating model performance concerning the growing hierarchical complexity in cells with critical information becomes paramount. To tackle this, we annotate each example in Multihiertt with the hierarchy depth of cells containing relevant information. For table with multiple relevant cells, we consider the cell with the highest hierarchical depth for our analysis. Our approach to estimating hierarchy depth is illustrated in Figure 10. Figure 7(a)(a) in Appendix A.4 illustrate how we calculate hierarchical complexity for examples with multiple relevant rows at various hierarchical depths.
5. Missing Information:
@@
Assessing empty cell proportions is crucial to quantify data ambiguity. More empty cells suggest higher uncertainty, which can hinder models’ ability to derive meaningful insights and impact reasoning accuracy. In Multihiertt, where tables are hierarchical in nature and empty cells occur quite frequently, we annotate examples with the empty cells percentage, contributing to our understanding of data ambiguity. For distribution of missing information (empty cells proportions) across datasets, refer to Figure 7(a) (b) in Appendix A.4.
Annotation Splits.
[…]
3 Experimental Results
@@
In this study, we choose to experiment with LLMs such as GPT-3.5-Turbo, GPT-4, PaLM-540B, Mistral-7B-Instruct [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, Llama-2-13B [^3] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf and MAmmoTH-13B [^4] https://huggingface.co/TIGER-Lab/MAmmoTH-13B. These LLMs are at the cutting edge for both open-source and closed models applications. Models like MAmmoTH-13B are specifically fine-tuned during pre-training to excel in mathematical reasoning tasks. For more detail about the the model choices refer to Appendix A.2.
LLMs Prompting Methods:
@@
For an instruction-tuned LLM, it’s assumed that we give the model a task-specific instruction $\mathcal{I}$ accompanied with a few (usually $k\in\{2,4\}$) demonstrations $\mathcal{D_{T}}$ of a task T. We experiment with the following prompting techniques:
@@
(a.) Direct: In this setup, we explicitly instruct the models to abstain from providing explanations and just return the final answer to the posed question. For this scenario, $\mathcal{D_{T}}$ contains $\{(p_{i},q_{i},a_{i})\}_{i=1}^{k}$ where p is the premise (table-text), q is the question, and a is the ground-truth answer.
@@
(b.) CoT: Moving forward, we experiment with the chain-of-thoughts prompting strategy where we instruct the model to output the explanation to the answer derived by it. Here, our $\mathcal{D_{T}}$ contains $\{(p_{i},q_{i},e_{i})\}_{i=1}^{k}$ where p is the premise which includes the table and the associated text, q is the question and e is the explanation of the answer.
@@
(c.) PoT: In this case, the expected response is a code derivation of the answer. Here, $\mathcal{D_{T}}$ contains $\{(p_{i},q_{i},c_{i})\}_{i=1}^{k}$ where p is the premise which includes the table and the associated text, q is the question and c is the code-derivation of the answer.
@@
(d.) Decomposers: Ye et al. (2023) proposed to address the challenge of handling large tables by decomposing them into more manageable subtables. Similarly, complex questions are handled by breaking them down into simpler subquestions. Decomposition proves effective with SQL tables, facilitating the removal of distracting details while retaining all supporting evidence. Questions are first parsed to break them down into simpler, more manageable subquestions. The model then addresses each subquestion independently before composing the answers to arrive at the final solution. In this case, our demonstration set $\mathcal{D_{T}}$ contains $\{(p_{i}^{{}^{\prime}},\langle q_{1},q_{2},...,q_{n}\rangle,a_{i})\}_{i=1}^{k}$ where $p^{{}^{\prime}}$ is the premise obtained by the irrelevant information removal to the question from the original premise p and $\langle q_{1},q_{2},...,q_{n}\rangle$ are the subquestions whose answers lead to the final answer.
EEDP Prompting Strategy:
@@
We propose a novel prompting strategy: Elicit $\longrightarrow$ Extract $\longrightarrow$ Decompose $\longrightarrow$ Predict. Figure 4 show an illustration of our EEDP approach. Below are the detail of each EEDP step:
@@
    Predict: The model finally returns the derived answer in the above steps.
Figure 11 shows a example for EEDP strategy with one shot.
Dataset Model Direct CoT PoT EEDP Decomposers
GPT-4 55.81 86.91 89.99 88.67 47.46
GPT-3.5-Turbo 31.38 77.57 82.11 79.73 28.53
TATQA PaLM 2-540B 44.66 62.93 61.60 81.51 57.94
Llama 2-13B 3.36 35.95 34.16 40.95 25.93
MAmmoTH-13B 19.11 56.25 10.02 4.37 22.89
Mistral-7B 10.92 59.14 16.53 56.06 7.24
GPT-4 65.12 72.38 75.26 76.05 44.93
GPT-3.5-Turbo 40.47 59.18 68.97 61.88 32.33
FinQA PaLM 2-540B 30.33 34.79 30.41 61.95 46.38
Llama 2-13B 1.80 25.34 12.97 30.47 11.91
MAmmoTH-13B 22.83 35.32 15.86 35.05 17.65
Mistral-7B 26.11 34.23 10.56 34.86 12.34
GPT-4 63.10 71.19 78.81 77.91 18.76
GPT-3.5-Turbo 37.62 48.33 61.19 61.75 10.50
ConvFinQA PaLM 2-540B 20.19 38.00 40.14 63.42 22.32
Llama 2-13B 3.80 29.45 29.92 39.42 10.35
MAmmoTH-13B 21.61 46.08 8.78 32.77 7.83
Mistral-7B 12.35 48.45 14.48 36.57 11.16
GPT-4 41.35 55.13 67.23 70.32 36.86
GPT-3.5-Turbo 25.88 42.33 52.18 49.65 20.61
Multihiertt PaLM 2-540B 14.20 20.67 36.52 37.97 20.19
Llama 2-13B 1.54 30.66 18.12 24.15 16.86
MAmmoTH-13B 10.12 18.56 6.57 18.36 11.87
Mistral-7B 14.909 22.92 14.94 10.97 11.63
@@
Results and Analysis.
Table 2 shows a comparison in performance between different prompting strategies. Despite being a single prompt, EEDP demonstrates comparable or superior performance compared to PoT. Notably, we outperform PoT significantly for PaLM-2-540B and LLAMA-2-13B across most datasets. Moreover, while PoT relies on external tools for executing mathematical programs/code to obtain answers, EEDP exclusively utilizes LLM for all tasks, including evidence extraction, operation identification, and execution, ensuring precision throughout the process.
As shown in Table 2, the Decomposers prompting strategy exhibits a much poorer performance compared to other strategies. The reason behind this was statistically found to be the inaccurate formation of subtables from the main table, leading to information loss as described in the previous paragraph. The performance of EEDP either surpasses or matches very closely with that of PoT. The number of shots was adjusted depending on the context length of the model.
We can see that MAmmoTH-13B model, which is fine-tuned on the MathInstruct dataset Yue et al. (2024) containing Instruction-Response pairs where the responses are a hybrid of CoT and PoT rationales, fails to perform well with the EEDP methodology. We argue that this is due to two potential reasons: (a.) Reduction of the number of shots to adjust the context length as the EEDP response is longer than that of the other methods, and (b.) Finetuning may contribute to suboptimal performance due to its alignment with a particular style and format of responses, potentially limiting the model’s adaptability and generalization to other diverse contexts.
EEDP’s Computational Efficiency
EEDP functions as a unified single-prompt method, minimizing computational complexity. Unlike methods like PoT, which rely on external tools, EEDP operates independently. When assessing computational cost, we consider API calls and token generation. Since EEDP uses a single-step prompting approach, only one API call is needed per query, making its computational cost comparable to methods like CoT. For inference with open-source models, we used hardware with an A40 40GB GPU. Processing one dataset per model using the vLLM library took approximately 10 hours.
4 Where do LLMs fail?
Through manual inspection, we rigorously evaluate the EEDP responses against the meta-annotations from section 2 as ground-truth benchmarks for extraction and model reasoning accuracy. The reasoning programs represent sequences of arithmetic operations necessary to derive the final answer, utilizing values extracted from supporting evidence as operands. To assess calculation accuracy, we manually identify the model’s instantiation and precision errors. Our EEDP prompt ensures that the model predominantly outputs responses in the expected format, with exceptions being rare. However, since we manually analyze all outputs, we do not penalize the model for format deviations but rather for incorrect outputs. Penalties are applied only when the model makes errors in extraction, reasoning, and/or calculation. Below, we categorize the EEDP response errors in detail based on their origins:
Dataset Error Type Per.(%)
FinQA Extraction E1 10.38
E2 25
Reasoning R1 25
R2 15.57
Calculation C1/C2 24.06
ConvFinQA Extraction E1 8.45
E2 14.08
Reasoning R1 19.72
R2 36.62
Calculation C1/C2 21.13
TATQA Extraction E1 13.79
E2 31.03
Reasoning R1 22.41
R2 5.17
Calculation C1/C2 27.59
Multihiertt Extraction E1 20.5
E2 31.5
Reasoning R1 15.5
R2 12
Calculation C1/C2 20.5
Table 3: Error Analysis on Various Datasets. In this table, Extraction.E1: Missing Evidences, Extraction.E2: Wrong Evidences, Reasoning.R1: Insufficient Domain Knowledge, Reasoning.R2: Question Misinterpretation, Calculation: Instantiation (C1) and Precision errors (C2)